OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality x - 3 < -11 or x + 5 greater than or equal to 14 A number line with a closed circle on -8, shading to the left, and an open circle on 9, shading to the right. A number line with an open circle on -8, a closed circle on 9, and shading in between. A number line with an open circle on -8, shading to the left, and a closed circle on 9, shading to the right. A number line with a closed circle on -8, an open circle on 9, and shading in between.

OpenStudy (anonymous):

@.Sam. please help sam

OpenStudy (anonymous):

@Turtledove

OpenStudy (anonymous):

nice to see you again

OpenStudy (texaschic101):

x - 3 < -11 x + 5 >= 14 x < -11 + 3 x >= 14 - 5 x < - 8 x >= 9 3rd answer choice is correct :)

OpenStudy (texaschic101):

hello to you too :)

OpenStudy (anonymous):

ty

OpenStudy (texaschic101):

anytime.....glad I could help:)